### v2

# Opening Articulated Objects in the Real World

Abstract
What does it take to build mobile manipulation systems that can competently operate on previously unseen objects in previously unseen environments? This work answers this question using opening of articulated objects as a mobile manipulation testbed. Specifically, our focus is on the end-to-end performance on this task without any privileged information, i.e. the robot starts at a location with the novel target articulated object in view, and has to approach the object and successfully open it. We first develop a system for this task, and then conduct 100+ end-to-end system tests across 13 real world test sites. Our large-scale study reveals a number of surprising findings: a) modular systems outperform end-to-end learned systems for this task, even when the end-to-end learned systems are trained on 1000+ demonstrations, b) perception, and not precise end-effector control, is the primary bottleneck to task success, and c) state-of-the-art articulation parameter estimation models developed in isolation struggle when faced with robot-centric viewpoints. Overall, our findings highlight the limitations of developing components of the pipeline in isolation and underscore the need for system-level research, providing a pragmatic roadmap for building generalizable mobile manipulation systems. Videos, code, and models are available on the project website: https://arjung128.github.io/opening-articulated-objects/.
I Introduction
Developing mobile manipulators that can reliably perform everyday tasks in diverse environments remains a fundamental challenge in robotics. A major obstacle to realizing this vision lies in the lack of strong generalization capabilities: current systems struggle to adapt to novel objects and unfamiliar settings. This difficulty stems from the inherent complexity of mobile manipulation in arbitrary environments, which requires robust perception, precise motion planning, and successful execution. While considerable efforts have been devoted to improving each of these subproblems in isolation, relatively few studies bring these components together into a cohesive system designed for real world deployment. Building and studying end-to-end systems rather than isolated subproblems can reveal hidden challenges crucial for practical deployment. By focusing on the full system, research can shift attention towards overall task success rather than marginal gains in individual components that may not be the primary bottleneck. Therefore, prioritizing system-level end-to-end performance—particularly the ability to generalize to unseen objects in unfamiliar settings—is essential for achieving reliable and capable mobile manipulators in everyday environments.
To tackle the challenge of strong generalization within the context of mobile manipulation, we work on the task of opening articulated objects in diverse previously unseen real world environments (Figure LABEL:fig:dummy-teaser). Articulated objects, such as cabinets, drawers, and ovens, are ubiquitous in indoor environments, making their reliable manipulation a key capability for robots operating in the wild. In our problem setting, a Stretch RE2 robot (a commodity mobile manipulator without any hardware modifications) is placed in a novel environment in front of a previously unseen articulated object, and the objective is to solve the full end-to-end task. This includes detecting the object and estimating its articulation parameters, generating a precise whole-body motion plan, navigating to an optimal interaction position, securing a firm grasp on the handle, and successfully opening the object without collisions—all performed zero-shot, without any privileged information about the object or environment. Thus, this problem setting includes all of the key challenges that a mobile manipulator would face when interacting with previously unseen objects in unfamiliar environments in the real world, and thus forms a good test bed for study. We first developed an end-to-end mobile manipulation system to tackle this task and then conducted extensive end-to-end testing to understand the current bottlenecks in building such a system. This paper describes the system we build, the testing we conducted, and the lessons we learned along the way.
The System. We considered two broads ways of putting together such a system: a modular approach and an end-to-end learning approach, but ultimately favored a modular approach. Our approach, called MOSART for a MOdular System for opening ARTiculated objects, develops and utilizes state-of-the-art modules for perception, planning, and adaptation (shown in Figure I). Very briefly, MOSART adapts a Mask RCNN model [23] for inferring articulation parameters, extends a trajectory optimization framework for producing whole body motion plans [21], and utilizes proprioceptive feedback for mitigating last centimeter errors.
The Testing. We conduct large-scale tests to assess the capabilities of MOSART. This testing is conducted in 13 test sites from 10 different buildings across 31 different articulated objects in the wild. Testing sites include offices, classrooms, apartments, office kitchenettes, and lounges (see Figure LABEL:fig:dummy-teaser). None of these test objects or sites were used for development in any way, and testing was done just once to mitigate any influence on the design of the system. Each trial started with the robot being placed such that the object was in view. A user selected which articulated object was to be opened and hit the go button. If the drawer was opened more than 24cm or a cabinet was pulled open by $60^{\circ}$, the trial was deemed a success. Finally, we also conduct experiments to understand a) how MOSART compares to an end-to-end learning approach, b) how sensitive MOSART is to the performance of each individual submodule, c) whether MOSART can generalize to diverse handles and other articulation types, and d) what are the biggest bottlenecks that cause MOSART to fail, providing insights into open challenges in building generalizable mobile manipulation systems.
Figure 0: MOSART Design. The perception module outputs 3D articulation parameters in the robot frame using RGB-D images. The robot then navigates to the target location based on articulation type. Next, we use SeqIK to find a whole-body motion plan. We execute the first robot configuration from the plan to obtain a pre-grasp pose. We then use our contact-based adaptation for improved grasping. Once the handle is grasped, we execute the rest of the plan.
The Findings. Our large-scale system-level study revealed various takeaways:
• A modular system outperforms an end-to-end learning system when tested for generalization to unseen objects in novel environments. We compare MOSART to the recent and contemporary imitation learning work Robot Utility Models (RUM) from Etukuru et al. [12]. RUM is trained on a large-scale dataset of expert demonstrations: 1200 demos for opening cabinets, and 525 demonstrations for opening drawers across approximately 40 environments, by far the largest imitation learning dataset for this task. To our surprise, we find that the modular system outperforms this latest end-to-end learning method (Section IV-B and Table I). This results is particularly useful in context of the ongoing debate between modular learning and end-to-end learning in robotics at large, and specifically for recent end-to-end learning works that tackle mobile manipulation problems [53, 63, 65, 20].
• State-of-the-art perception modules for articulation parameter prediction struggle on images from a robot. Significant progress has been made in predicting articulation parameters from images [57, 29, 7, 48, 40, 61, 45, 41]. However, most of this progress has been made in isolation, without evaluating these methods as part of a complete, end-to-end system on real-world, in-the-wild images. When tested on viewpoints from the robot in the real world, recent state-of-the-art approaches such as OPD-Multi [57] and AO-Grasp [41] struggle to make accurate predictions. In contrast, our choice to develop a Articulation-parameter Prediction Module (APM) targeted towards images that the robot was expected to see, achieves substantially better performance vs. OPD-Multi [57] (Section IV-C1 and Figure 4) and AO-Grasp [41] (Section IV-C1, Table III, Figure A16).
• Modular systems can generalize to diverse handles and be easily extended to other articulation types. Not only did the modular system MOSART outperform the competing imitation learning method RUM [12], we found the modular system to be advantageous in other ways. MOSART is able to generalize to diverse handles (Section IV-E2). As vision foundation models keep improving, a modular system continues to automatically improve. Additionally, we found it easy to extend our modular system to a new articulation type (horizontal-hinged toaster ovens, Section A10). In comparison, an imitation learning system will need to recollect a large amount of training data for tackling a new articulation type.
• The failure mode analysis reveals that robust perception, and not control, is the biggest bottleneck in overall task success. Surprisingly, perception and not the precise control of the end-effector (to follow the handle as the object articulates) is what makes this problem hard, contrary to what prior work focuses on [21, 30, 9]. Inability to detect objects and handles accounts for 59% of the failures of the system (Figure 6, Section IV-F). Furthermore, the specific ways in which the perception system failed was also revealing. It is not as much a failure in estimating articulation parameters, but the detection of target objects and estimation of the handle location in 3D are the bottlenecks. Specifically, cabinets with meshed surfaces do not get detected and keyholes get mistaken as handles. It may be important to capture data for such corner cases.
• Last-centimeter grasping remains challenging. Control was surprisingly robust to mis-estimations in the articulation parameters. Once the end-effector acquired a firm grasp of the handle, we found that the system is able to sufficiently open the cabinet even when the radius is off by as much as $10cm$ (Section IV-D). However, securing this initial grasp proved to be somewhat challenging, with last-centimeter grasping errors accounting for 25% of all failures. While proprioceptive feedback provided a simple yet effective adaptation mechanism, it did not fully eliminate last-centimeter grasping errors. Tackling such last-centimeter errors remains an avenue for future work. Some form of a closed-loop visual grasping controller that is robust to slight mis-calibration and imprecise navigation may improve performance on mobile manipulation tasks which require precise grasping.
Broader Implications. Our findings provide key insights for designing generalizable mobile manipulation systems, offering several takeaways that should inform future research and system development. In particular, our large-scale, real-world experiments show that vision models trained in isolation frequently underperform when confronted with the distinct viewpoints encountered by robots, highlighting the need for comprehensive, system-level research. Moreover, we identify perception as the prsincipal bottleneck in achieving consistent task success, a finding that underscores the need for future research to prioritize robust perception modules. Additionally, our results indicate that modular approaches tend to outperform end-to-end methods, cautioning practitioners that scaling imitation-learning datasets to even 1,000 demonstrations does not necessarily lead to broad generalization. Notably, modular systems also facilitate easier adaptation to related tasks without requiring extensive new data collection. Finally, last-centimeter grasping remains a key challenge. While contact correction helps alleviate this failure mode to some extent, developing closed-loop interaction strategies could improve performance in tasks requiring precision, making it an important direction for future research. Overall, these lessons provide a pragmatic roadmap for researchers and practitioners aiming to build generalizable mobile manipulation systems.
Figure 1: Overview of the Articulation-parameter Prediction Module (APM). Given an RGB image our modified Mask RCNN detects articulated objects and predicts the articulation type, the handle orientation, the 2D segmentation mask, and the 2D handle keypoint. We fit a convex hull to the segmentation mask and simplify it to a quadrilateral. We fit a plane to the depth image points that lie inside the mask to estimate the surface normal. The 2D handle and quadrilateral corners are lifted to 3D using the depth image. All predictions are transformed to the robot base frame. The final output of the module includes the 3D handle coordinate and surface normal in the base coordinate frame for all articulated objects, and additionally the radius and rotation axis for cabinets.
II Related Work
II-A Predicting Articulation Parameters
Researchers have looked at different aspects: a) construction of datasets (from simulation [33, 10, 16], real world images [57, 29, 7], and real world 3D scans [21, 58]), b) use of different input modalities to make predictions (RGB images [57], point clouds [36, 44, 60, 11, 69, 35, 51], RGB-D images [29, 64, 1, 68]), c) use of diverse sources of supervision [48], and d) predicting sites for interaction (i.e. handles) in addition to articulation parameters [40, 61, 45, 67]. Most past works are largely evaluated in simulation, and because no prediction is made for other articulation parameters (e.g. radius), these works cannot be used to generate motion plans to fully articulate objects. Sun et al. [57] add additional heads to Mask RCNN to predict articulation parameters in 3D. We also add additional heads to Mask RCNN but rather than directly predicting 3D outputs using RGB-D input, we take a 2D prediction from RGB and 3D lifting using depth image approach [22, 62].
II-B Generating Motion Plans
Opening doors and drawers requires the end-effector to conform to the constraint defined by the object handle’s trajectory (which in turn is determined by the location of the articulation joint and the handle). Depending on the robot morphology, robot configurations that conform to this end-effector constraint might represent a measure zero set. This makes it challenging to directly use sampling-based motion planners [31, 34]. Past research has therefore developed specialized methods for planning under such constraints [32, 4] and used it to articulate objects [5, 9, 50, 47, 38, 6, 59, 24, 25, 42]. Another line of work casts it as trajectory optimization [21, 71, 52] or optimal control  [13, 46, 54, 39] problem. As these only search for a solution locally (vs. motion planning that searches globally via sampling), it is important to properly initialize the trajectory optimizers. Recent work from Gupta et al. [21] alleviates this limitation by using learning to predict good initializations for trajectory optimization thus generating high-quality motion plans quickly. We adopt their approach but extend it to produce whole body motion plans as we describe in Section III.
II-C Mobile Manipulation
Recent papers have looked at different aspects: pick-move-place tasks [66], high-level planning given natural language instructions [2], dynamic whole body control [14], building simulators [43], developing tele-operation setups [15]. [3] tackle a grocery shopping scenario with a custom robot, and conduct extensive field tests over 18 months. Many recent works adopt an end-to-end learning approach for opening articulated objects [65, 53, 63, 26, 12]. [65] also work with a Stretch Robot but focus on sim2real transfer and modify the environment to simplify the challenges due to closed-kinematic chains. The Dobb-E system from [53] showcases interactions with articulated objects, but requires retraining on each test object using human-collected demonstrations. Work from [63] replaces the need for test-time demonstrations with test-time adaptation via RL requiring about an hour of interaction for the adaptation. [55] enable a quadruped to articulate heavy doors and dishwashers, but assume privileged environment information (e.g. a model for the door). Recently, Etukuru et al. [12] introduced RUM, a large-scale imitation learning system for opening articulated objects, collecting over 1,000 demonstrations across approximately 40 environments. However, their approach relies on ground-truth articulation type, an eye-in-hand camera, an approximate handle height, and requires the robot to be positioned directly in front of the object with a clear view of the handle. In contrast to these approaches, we develop a system that operates on novel object instances in novel environments in a zero-shot manner without requiring any privileged information.
Figure 2: Topdown Navigation Targets and Corrective Motions. We show the topdown navigation targets relative to the handle for each articulation type. For left-hinged cabinets, correction is a rotation in $1^{\circ}$ increments. For the other objects, we extend the arm in $1cm$ increments.
III Modular System for Opening Articulated Objects (MOSART)
We follow a modular approach comprising of a) a Perception Module that predicts articulation parameters using on-board an RGB-D camera, b) a Motion Plan Generator that converts predicted articulations into whole-body motion plans, and c) an Execution Module that adapts and executes the generated motion plan using proprioceptive feedback. Figure I shows an overview. We describe the pipeline in terms of two prototypical articulated objects: drawers for prismatic joints and cabinets for hinged joints, but our pipeline is more general as our experiments will reveal.
III-A Predicting articulation parameters using on-board RGB-D images (APM)
Our Articulation-parameter Prediction Module (APM), shown in Figure 1, detects and predicts articulation parameters for cabinets and drawers from RGB-D images. These articulation parameters include the 3D handle location and surface normal for drawers, and additionally the rotation axis and radius for cabinets. We predict 2D quantities from RGB images, and lift them to 3D using the depth image.
For 2D prediction from RGB images, we adopt Mask RCNN [23]. As is, Mask RCNN predicts a 2D segmentation mask and the class of each detected object (in our case, the articulation type: drawer, left-hinged cabinet, or right-hinged cabinet). We add additional heads to Mask RCNN to predict the handle’s 2D pixel location and orientation (horizontal or vertical). Both additional heads are treated as classification tasks and trained with a cross-entropy loss. For the 2D handle coordinate prediction, we minimize the cross-entropy loss over a 2D spatial map.
We use the depth image to lift these 2D predictions to 3D. For the surface normal, we fit a plane to the 3D points within the predicted segmentation mask. For the 3D handle position, we lookup the 3D coordinate corresponding to the predicted 2D handle coordinate. For cabinets, we also need the radius and the axis of rotation. We compute the convex hull of the predicted 2D segmentation mask, and simplify it to a quadrilateral. We lift the vertices of this quadrilateral to 3D and infer the rotation axis from the corners, e.g. for a left-hinged cabinet we define the rotation axis using the left corners. We use the distance of the handle to its projection on the rotation axis as the radius.
We train our modified Mask RCNN on the ArtObjSim dataset [21]. ArtObjSim contains 3500+ articulated objects across 97 scenes from the HM3D dataset [49]. Each articulated object comes with 3D annotations for its extent, handle location, articulation type, and articulation parameters. We train on images and 2D annotations rendered out from arbitrary locations in the scene.
III-B Motion plan generation
Given waypoints computed using our predicted articulation parameters, our motion plan generator synthesizes a navigation target and a whole-body motion plan to open the given articulated object in a collision-free manner. These waypoints, representing end-effector poses, are determined based on the predicted articulation type and handle location: a linear trajectory along the predicted surface normal for drawers, and a quarter-circle path also using the predicted radius for cabinets. We build upon past work [21] that converts ground-truth end-effector pose trajectories into robot joint angle trajectories. Specifically, rather than casting it as a constrained motion planning problem, Gupta et al. [21] view it as a trajectory optimization problem and design SeqIK, a trajectory optimizer specifically suited to this task. SeqIK translates an initial robot configuration (base position and arm joint angles denoted by $\theta_{0}$) into a strategy that can be decoded into a motion plan (desired joint angle trajectory) when provided with a desired end-effector pose trajectory ${\bf w}$, via $\text{SeqIK}(\theta_{0})({\bf w})$. SeqIK performs inverse kinematics calls sequentially, warm-starting the next inverse kinematics call with the output of the current. This leads to accurate motion plans with only a few IK calls.
Figure 3: Contact Correction. (a) shows a grasping attempt with No contact correction, whereas (b) shows the grasping attempt with contact correction. Due to compounding errors and shrinkage of the gripper when closing, the without contact correction version fails to grasp, whereas our contact-based correction mechanism leads to a successful grasp.
We extend the framework from [21] in three ways. First, [21] works with the Franka Emika Panda robot. We adopt their implementation to work with the Stretch RE2 robot, which has fewer degrees of freedom. Second, we work with articulation parameters predicted using APM as opposed to ground truth articulation parameters. Finally, we obtain whole-body motion plans rather than base placement + arm-only plans developed in [21]. We find whole-body planning essential to fully open a wide variety of cabinets and drawers due to the limited number of degrees of freedom of the Stretch RE2. Here, in each SeqIK step, given the fixed initial base pose, we allow inverse kinematics to search for base rotation in addition to the arm joints.
SeqIK requires an initial base position and arm joint angles $\theta_{0}$. For the initial base position, we utilize MPAO (No neural network), the data-driven method from [21] that ranks robot configurations (base positions and arm joint angles) by how successfully SeqIK can decode these configurations into collision-free constraint-abiding motion plans. Figure 2(a) shows the base positions found by this procedure for drawers, left-hinged and right-hinged cabinets. We use these as the navigation targets for each articulation type respectively. For the initial arm joint angles $\theta_{0}$, we find that using the default neutral joint angles works well for the Stretch RE2.
III-C Adapting and executing motion plans using proprioceptive feedback
Minor errors in state estimation, navigation, and calibration compound to prevent handle grasping. Specifically, there is little tolerance in the depth direction. This is further exacerbated by the hardware design of the gripper that becomes shorter upon closure. We combat this via a contact-based correction strategy for arriving at the pre-grasp pose: we extend the gripper towards the object until contact is made. See Figure 2 for a visualization of the strategy for the different articulated objects. As is, the end-effector approaches the handles from the side, which works well for horizontal handles, but poses issues for vertical handles. For objects with vertical handles, we thus have an additional vertical primitive before the contact-based correction mechanism. We use the arm effort signal and the gripper’s yaw effort signal to detect contact. The change in the end-effector pose during contact correction phase provides feedback to update the desired trajectory and motion plan. We ablate the use of contact correction in Section IV-C2.
Drawer Left-Hinged Right-Hinged Total
RUM [12] 2/3 1/3 0/3 3/9
MOSART (Ours) 3/3 2/3 2/3 7/9
Table I: Comparison of MOSART vs. RUM [12], a recent large-scale end-to-end imitation learning method trained on 1200 demos for opening cabinets and 525 demos for opening drawers across 40 different environments. Our evaluation spans objects in-the-wild from four previously unseen environments across three buildings.
III-D Full end-to-end execution
Starting from where the robot can see the target articulated object, we use APM to predict the handle location (in the camera’s coordinate frame), the surface normal, and the radius (for cabinets). We use a calibrated robot URDF to transform these 3D predictions from the camera frame to the base frame. Next, we generate a whole-body motion plan and execute the first qpose (full robot configuration), followed by our contact-based correction mechanism. Once the handle has been grasped, we execute the rest of the motion plan.
IV Experiments
We work with the Stretch RE2 robot. We first present our end-to-end system test results, evaluating MOSART on 31 novel drawers and cupboards across 10 buildings (Section IV-A). To see how a modular system compares to an end-to-end learning approach, we compare MOSART to RUM [12] and a sim2real imitation learning approach (Section IV-B). To further understand the system, we then evaluate individual pipeline modules (Section IV-C). This includes evaluating the quality of our MaskRCNN-based perception module (as well as a Detic-based perception model) on real world images, comparing APM to two recent articulation parameter prediction systems [41, 57], and an ablation for the contact-based correction mechanism. To understand the interaction between perception and control, we quantify the robustness of execution to state estimation inaccuracies (Section IV-D). We then study the generalization of our pipeline to other articulation types and diverse handles (Section IV-E), before we analyzing the failure modes of our system (Section IV-F).
IV-A End-to-end System Tests
We test our end-to-end system across 8 office buildings and 2 apartments on a total of 31 distinct articulated objects. These test objects do not overlap with ones used for development. In each test, the robot is placed approximately $1.5m$ from the target object with the camera oriented so as to have the target object in view. We introduce variation in the starting pose of the robot to test the robustness of the approach but use the same starting pose when comparing different methods. A trial is successful if the drawer is pulled out by $24cm$ / the cupboard is opened more than $60^{\circ}$.
Results. Overall, our system achieves a 61% success rate across 31 unseen cabinets and drawers in unseen real world environments. Figure 1 shows examples of deployments of our full pipeline. For most successful trials, the robot opens the drawer / cupboard completely (i.e. drawers by $35cm$ and cupboards by $90^{\circ}$) in a graceful manner (see videos in supplementary materials). Section IV-F provides a extensive discussion of the failure modes.
Table II: Accuracy of APM predictions on images collected during our large-scale real world evaluation.
APM with Mask RCNN APM with Detic
Drawer Left-Hinged Right-Hinged All Drawer Left-Hinged Right-Hinged All
Detection 9/9 8/9 12/13 29/31 9/9 9/9 13/13 31/31
Handle orientation 9/9 8/8 11/12 28/29 9/9 9/9 12/13 30/31
Articulation type 9/9 8/8 12/12 29/29 9/9 9/9 12/13 30/31
Mean handle error 3.17 cm 1.43 cm 1.35 cm 2.11 cm 1.51 cm 0.82 cm 1.34 cm 1.27 cm
Mean radius error n/a 0.47 cm 1.32 cm 0.98 cm n/a 3.28 cm 2.31 cm 2.70 cm
IV-B Comparisons to End-to-End Imitation Learning
IV-B1 Robot Utility Models (RUM)
Next, we investigate how MOSART compares to a large-scale end-to-end imitation learning approach in generalizing to novel objects in previously unseen environments. Specifically, we evaluate MOSART against Robot Utility Models (RUM), recently introduced by Etukuru et al. [12]. RUM is trained on an extensive dataset of expert demonstrations: 1,200 for opening cabinets and 525 for opening drawers, collected across approximately 40 environments. This dataset is by far the largest imitation learning dataset for articulated object manipulation, making RUM a strong baseline for our comparison.
Unlike MOSART, RUM operates under several assumptions. It requires prior knowledge of the articulation type (e.g. drawer vs. cabinet), an approximate handle height, and it assumes the robot is positioned directly in front of the object with an optimal view of the handle. To overcome these assumptions, we leverage MOSART’s APM to infer the articulation type and handle height, and we use MOSART’s navigation module to position the robot directly in front of the object (centered on the handle), among other minor adjustments described in Section A11-A.
As shown in Table I, MOSART substantially outperforms RUM in a paired evaluation on unseen objects across four novel environments. One common failure mode we observe for RUM is the inability to grasp the handle of the object, despite getting very close to it. Another is incorrect handle selection, particularly when multiple handles are in close proximity and within view. This stems from RUM’s inability to explicitly specify the task, making it susceptible to such ambiguities and highlighting a key limitation of the approach. Furthermore, our evaluation focuses solely on in-the-wild objects, where such challenging cases naturally arise, whereas a considerable portion of RUM’s evaluation was conducted on isolated objects in controlled lab environments. These findings highlight a) the importance of end-to-end system tests in-the-wild, b) the challenging nature of this problem that requires generalization to previously unseen objects in novel environments, and c) the effectiveness of MOSART.
Drawer Left-Hinged Right-Hinged Total
AO-Grasp [41] 13.11 cm 33.66 cm 24.46 cm 23.36 cm
Ours 3.17 cm 1.43 cm 1.35 cm 2.11 cm
Table III: Quantitative Comparison to AO-Grasp. Mean handle error across all instances of our real world testing.
IV-B2 Sim2Real Behavior Cloning
We also compare to a Sim2Real behavior cloning approach, building on top of prior work [21]. Here, we start with their open-loop imitation learning method and train it on the articulated objects in ArtObjSim [21]. The policy consumes an RGB-D image and outputs whole-body motion plans. Similar to [21], we find that in simulation, the learned model outputs plans that coarsely capture the overall motion but aren’t precise enough. We also deploy predicted plans onto the Stretch in the real world, but the model struggles further due to the sim2real gap, achieving a 0% success rate.
We try to improve this sim2real policy by fine-tuning it on real world data. We collect 32 demonstrations using tele-operation for each object used for developing our pipeline (one drawer and two cabinets), and learn a separate closed-loop behavior cloning policy for each object type. We initialize the visual backbone using the open-loop policy trained in simulation on ArtObjSim. We find that these policies only achieve a 70% success rate on the training objects in the real world, and completely fails on unseen objects. A diffusion policy [8] trained on this data also similarly struggled to generalize to novel objects in previously unseen environments.
Figure 4: Comparison to OPDMulti [57]. We perform a qualitative comparison to OPDMulti [57] (appeared in 3DV 2024) on the same six images presented in Figure A8. OPDMulti fails in various ways: a) segmentation masks bleed outside of the object, and b) merging of multiple objects into one. Our model produces more accurate segmentation masks (which affect the surface normal, and thus the navigation).
IV-C Evaluation of Individual Modules
IV-C1 Articulation Parameter Prediction Accuracy
We evaluate APM on all of the real world objects in our test set. We compute a) the detection accuracy, b) handle orientation accuracy, c) articulation type accuracy, d) mean handle location error, and e) mean radius error. The latter four are computed only for detected instances. We manually annotate to obtain 3D ground truth for evaluation.
We find that detection accuracy is high, but not 100%. APM fails to detect 2 of the 31 instances, see example missed instances in Figure 6. Articulation type prediction is always correct and the handle orientation is correct 28/29 times. Furthermore, the mean handle location error is small at $2.11cm$ across categories. The mean radius error is $0.98cm$ across categories, which contrary to our original belief, did not prove to be an obstacle during deployment. Table II breaks down results by categories and the video provides qualitative visualizations. The model detects fully visible objects well, and is sometimes also able to detect occluded objects.
We also compare APM to two recent works: AO-Grasp [41] and OPDMulti [57]. We perform a quantitative comparison to AO-Grasp. Table III reports a quantitative comparison of the mean handle error on our real world images (lower is better). Our method significantly outperforms AO-Grasp. See Section A12 for a qualitative comparison to AO-Grasp, and for more details. We also perform a qualitative comparison to OPDMulti (see Figure 4), and find that our model produces more accurate segmentation masks (which affect the surface normal, and thus the navigation). OPDMulti does not predict handle locations, disqualifying it as a replacement to APM, and preventing us from performing a quantitative comparison. Additionally, we also compare to an approach which directly predicts articulation parameters in 3D (Section A13).
Since our system is modular, and because we only require 2D predictions for handles and extent, we are able to replace the modified Mask RCNN with another perception model. In particular, we replace it with Detic, a foundation model for open-vocabulary detection [70]. While Detic can produce a segmentation mask for a given object and its handle, it doesn’t natively output the handle orientation. We design a simple decision rule based on the X and Y variance of the 3D points in the handle segment (in the base robot frame) to obtain the handle orientation. We find our Detic-based pipeline to work comparably to our Mask RCNN-based pipeline (see Table II).
Drawer Left-Hinged Right-Hinged Total
w/ contact correction 8/9 4/9 7/13 19/31
w/o contact correction 1/9 4/9 6/13 11/31
Table IV: Ablations. Success rates for our full pipeline with and without contact correction (across all objects) in the real world. Contact correction significantly improves the success rate.
IV-C2 Effectiveness of Adaptation Strategies
Contact Correction: As noted in Table IV inclusion of the contact correction adaptation strategy improves success rate from 35% to 61%. Figure 3 visualizes how contact correction aids grasping. All of this improvement comes from horizontal handles (primarily on drawers). We find contact correction to be a simple but effective strategy to tackle last-centimeter errors in grasping.
IV-D Evaluation of Interaction between Perception and Control
While the radius estimate from APM was off by $1cm$ on average, MOSART was still able to fully articulate the object in a majority of cases. A tight grasp provides a certain level of tolerance to inaccuracies in state estimates. Here, we systematically test this. We perturb the ground truth radius in increments of $2cm$, compute a motion plan with this inaccurate radius, and deploy this inaccurate plan. We use ground truth parameters for everything except the radius estimate to ensure a tight grasp of the handle. To our surprise, even with a $10cm$ error in the radius estimate, the robot is able to open the cabinet significantly. When the radius is overestimated by $10cm$, the cabinet opens to $75^{\circ}$, and even when underestimated by the same amount, it still reaches $55^{\circ}$, just short of our success threshold. This suggests that with a solid grasp of the handle, execution can succeed even with relatively inaccurate state estimates. Handle grasping may be more critical than millimeter-level accuracy in estimating articulation parameters.
IV-E Generalizing to Other Articulation Types and Handles
IV-E1 Other Articulation Types
While MOSART was developed on cupboards and drawers, we investigate whether it is general enough to handle other articulation types. Ovens and other bottom-hinged objects require a downward semi-circular motion to open. which is quite distinct from that of cabinets and drawers. We conduct experiments to study how well our full end-to-end pipeline (perception, navigation, and execution) fares for opening a toaster oven.
We evaluate our Detic-based pipeline on three novel toaster ovens in five previously unseen kitchens, and find that our pipeline achieves a 80% success rate (our Mask RCNN-based pipeline struggles to detect ovens due to the relative scarcity in the ArtObjSim dataset). In the trial that failed, Detic failed to detect the toaster oven.
IV-E2 Diverse Handles
We study whether our full end-to-end system can work on a diverse set of handles, varying in color, geometry, and material. To evaluate this, we install six diverse handles on unseen objects in our test set, as depicted in Figure 5. This is done for three objects in our test set, one of each articulation type (drawer, left-hinged cabinet, right-hinged cabinet), for a total of 18 trials. We follow the same protocol as our end-to-end testing in Section IV-A.
Our system is successful on all 18/18 trials. Some handles are more challenging than others, particularly ones which curve, due to reduced lateral tolerance during grasping. Nonetheless, we find that our system is able to successfully open unseen objects even with a diverse set of handles.
Figure 5: Diverse Handles. We test MOSART on 6 diverse handles on 3 test objects. MOSART succeeds on all 18 trials.
IV-F Failure Mode Analysis
While MOSART is able to solve a majority of the novel test objects, our large scale evaluation reveals unforeseen failure in perception, navigation, and execution as summarized in Figure 6. 59% of failures (i.e. 7 failures) are due to perception, including various kinds of failures, such as failure to detect meshed cabinets (2/7), incorrect handle orientation prediction (1/7), bad radius prediction (2/7), and confusing keyholes for handles (2/7). These perception errors are due in part to testing on out of distribution objects. APM is trained on luxury homes from HM3D [49], whereas we mainly test on office buildings and apartments. 25% of failures are in execution wherein a firm centered grasp on the handle could not be acquired due to slight errors in robot calibration and navigation. The robot was able to start opening the object but eventually the handle slipped out of the gripper. The remaining 15% of failures are due to navigation errors on floors with thick carpets where the robot would audibly strain while rotating. The robot didn’t get into the predicted pre-grasp pose which led to a failed handle grasp.
The modular design of MOSART and our approach of making 3D predictions via lifting 2D predictions in APM, allows us to assess if use of foundation models (e.g. Detic [70], an open vocabulary object detector trained on broad data) can mitigate the errors due to perception (the largest failure mode). We find that Detic-based APM detects more instances than our Mask RCNN-based APM, but fails in other interesting ways. Thus, building stronger perception models in the context of robotics remains a major challenge in deploying such systems to the real world.
Figure 6: Failure Cases. Bar chart characterizing the various failure modes of MOSART for opening drawers and cabinets. 59% of failures are due to perception failures, including failure to detect the cabinets outlined and confusing keyholes for handles. Other failures were during execution, where the handle would slip out, and during navigation, where navigating on carpets was less accurate than on tiles.
V Limitations
While we evaluated MOSART on a variety of objects, our study excluded those with round knobs or no handles. MOSART uses pre-mined navigation targets without allowing for base translation, which may limit performance on objects surrounded by challenging collision geometry or with large-radii. Lastly, the mapping from articulation parameters to waypoints must be provided to MOSART, whereas an end-to-end learned approach may learn this from data.
VI Discussion
• Our large-scale experimental study reveals that the modular design of MOSART outperforms end-to-end imitation learning methods, even when the latter are trained on more than a thousand demonstrations (see Section IV-B). This finding can serve as guidance to practitioners, suggesting that scaling imitation-learning datasets beyond 1,000 demonstrations does not necessarily lead to generalization across diverse scenarios. While it is possible that imitation learning may perform better with even more data, scaling up such effort to collect even more data remains a challenge. MOSART’s modular design mitigates need for such data collection.
• State-of-the-art perceptions systems for inferring articulation parameters that have been developed as isolated modules [57, 41] do not work well on robot images on in-the-wild articulated objects. This finding underscores the need for holistic, system-level research to uncover previously overlooked bottlenecks that arise during real-world deployment. We develop APM, which does much better than prior work on previously unseen objects in novel environments from robot viewpoints in the real world.
• Modular systems offer distinct advantages over end-to-end learning approaches. As vision foundation models advance, a modular system inherently benefits from these improvements without requiring retraining. Moreover, we found that adapting our modular system to a new articulation type, such as horizontal-hinged toaster ovens, was straightforward. In contrast, an imitation learning system would require collecting additional training data to handle a new articulation type. These findings suggest that modular architectures not only enhance adaptability to new tasks but also enable systems to leverage ongoing advancements in perception without the overhead of continuous retraining.
• Perception remains the biggest bottleneck in deploying such a mobile manipulation system to in-the-wild settings. Perception failures (Mask RCNN, Detic) reveal the need for broader datasets spanning diverse environments (i.e. everyday cluttered offices / apartments beyond luxury homes). This highlights the importance of prioritizing advancements in perception modules to accelerate progress toward generalizable mobile manipulation systems.
• While the system demonstrated robustness to inaccuracies in articulation parameter estimation, successfully opening cabinets even with a 10cm error in radius (Section IV-D), achieving an initial secure grasp remained a critical challenge. Grasping failures accounted for 25% of all failures, highlighting the difficulty of precise last-centimeter grasping. Contact correction provided a partial remedy, but closed-loop visual grasping strategies of some form that can compensate for minor mis-calibrations and imprecise navigation could help further.
These insights offer valuable guidance for researchers and practitioners striving to develop mobile manipulation systems that generalize to unseen objects in real world environments.
Appendix
A7 Supporting Figures and Tables
Figure A7: Number of Waypoints Successfully Executed by MOSART. Histogram visualizing the number of waypoints executed, given a grasp of the handle is achieved. For most cases, all 10 waypoints are successfully executed and the object is fully opened.
Figure A8: Qualitative Results from APM. Output from APM’s Mask RCNN on images from our real world testing. We show the predicted segmentation masks in blue, the fitted quadrilaterals in red, the handle point in green, and the predicted articulation type as text on the object.
Figure A9: Corrective Motions. We visualize the corrective motions for the different articulation types. For left-hinged cabinets, this is a counter-clockwise rotation in $1^{\circ}$ increments. For right-hinged cabinets and drawers, we extend the arm in $1cm$ increments. See Figure 2(a) and Section III.
A8 Evaluation of Motion Plan Generation
Figure A10: Effectiveness of Whole-Body SeqIK. We plot success / failure status of whole-body SeqIK (Section III) for the cabinets in the ArtObjSim dataset as a function of radius and handle height. We observe that handles above a certain height are difficult due to hardware limitations of the Stretch RE2, and cabinets with a radius larger than a certain threshold are challenging due to our parameterization of motion plans.
Here, we study the effectiveness of our methodology for generating whole-body motion plans for opening objects in the ArtObjSim dataset using the Stretch RE2. Figure A10 plots the success of various cabinets in the dataset as a function of both the handle height and the radius. We notice two trends: cabinets which have handles higher than $1.2m$ and cabinets with radii larger than $0.5m$ cannot be solved. The former can be attributed to a hardware limitation of the Stretch RE2. The lift joint of the robot cannot achieve a height of $1.2m$ or higher, leading to this failure. The latter, on the other hand, can be attributed to our parameterization of motion plans. Our current formulation only allows for rotation of the base and no translation. Allowing for the translation of the base in addition to rotation while interacting with an object will likely allow the robot to obtain more complex motion plans, including the kind needed for large-radii cabinets.
A9 Detic [70]-based Articulation-parameter Prediction Module
Since our system is modular, and because we only require 2D predictions for handles and extent, we are able to replace the modified Mask RCNN with another perception model. We developed a version of APM that uses 2D predictions from Detic [70]. Detic is an open-vocabulary object detector trained on broad datasets with an Swin-B backbone [37]. Table II presents perception results with with Detic-based APM.
While Detic can produce a segmentation mask for a given object and its handle, it doesn’t natively output the handle orientation. We design a simple decision rule based on the X and Y variance of the 3D points in the handle segment (in the base robot frame) to obtain the handle orientation.
Detic detects more instances than Mask RCNN (Table II) and is about as accurate in predicting handle location and radius (means are slightly higher due to an outlier prediction but the medians are about the same).
Figure A11: Detic Failure. An example failure of using Detic to try and detect articulated objects. In this instance, Detic clumps multiple cabinets together
A10 Other Articulation Types
Figure A12: SeqIK Heatmap for Opening a Bottom-Hinged Object. Topdown heatmap depicting the success rate of various base locations. The object is placed at (0, 0). Darker red indicates more waypoints were achieved, where all ten waypoints are achieved in the darkest red region.
Figure A13: Perception and Execution for Opening a Toaster Oven. Left. The output of our modified Mask RCNN demonstrating the detection of the toaster oven. Right. At the end of execution of our full Mask RCNN-based end-to-end pipeline showing the toaster oven is fully open.
Even though our focus is opening cupboards and drawers, we also investigate whether our proposed system is general enough to handle other articulation types. Ovens and other bottom-hinged objects require a downward semi-circular motion to open. We conduct experiments to check: a) can SeqIK generate motion plans to open bottom-hinged objects with the Stretch, and b) how well does our full end-to-end pipeline (perception, navigation, and execution) fare for opening a toaster oven?
SeqIK to Generate Motion Plans for Bottom-Hinged Objects. We utilize SeqIK to mine motion plans for an object with a bottom-hinge. As with cabinets and drawers, we limit the motion plans to allow for base rotation and do not allow for base translation. We find that SeqIK is able to find a full motion plan for various different base initializations, as depicted in Figure A12.
Performance of End-to-End Pipeline. We re-train our modified Mask RCNN with the inclusion of ovens and run our entire pipeline, as is, for opening a toaster oven. Figure A13 shows the output of our modified Mask RCNN model, and the successful opening of a toaster oven using our full end-to-end pipeline.
Across five trials, our modified Mask RCNN-based pipeline achieves a success rate of 20%. In the trials which fail, our modified Mask RCNN either fails to detect the oven, or incorrectly predicts the handle to be close to the bottom edge of the detected object. This is somewhat expected as the ArtObjSim dataset contains only 226 bottom-hinged objects (ovens, dishwashers, etc.) vs. 1200 drawers and 1992 cabinets. The smaller training set leads to reduced perception performance on such objects.
Since our system is modular, and because we only require 2D predictions for handles and extent, we are able to replace the modified Mask RCNN with another perception model. In particular, we replace it with Detic, a foundation model for open-vocabulary detection [70]. While Detic can produce a segmentation mask for a given object and its handle, it doesn’t natively output the handle orientation. We design a simple decision rule based on the X and Y variance of the 3D points in the handle segment (in the base robot frame) to obtain the handle orientation. Across five trials, our Detic-based pipeline achieves a success rate of 80%, which is a lot higher than our modified Mask RCNN-based pipeline. This result is somewhat expected: Detic is a much bigger model with a stronger ViT backbone, trained on a much larger dataset than our modified Mask RCNN, which is why it seems to do better at detecting rarer objects.
A11 Implementation Details for Baselines
Here, we provide more details on comparisons to end-to-end learning baselines.
A11-A Robot Utility Models
We provide additional details about Robot Utility Models (RUM) [12]. Our task doesn’t provide any privileged information about the object or environment. RUM makes several assumptions, which we relax in the following ways to enable a comparison in our problem setting:
  • Articulation type. Firstly, RUM assumes the articulation type as input to the system. For this, we make use of our articulation type prediction from APM.
  • Approximate handle height. Next, RUM requires an approximate height of the handle to give the eye-in-hand camera an optimal view of the handle. For this as well, we make use of predictions of the 3D handle height from APM. To further benefit RUM, we try out two more heights in the vicinity, and report the best result for RUM.
  • Base location. Finally, RUM assumes the base of the robot is directly positioned in front of the articulated object. To relax this assumption, we allow the robot to navigate to one of two base locations: our pre-mined navigation target, or an ‘ideal’ location which is laterally centered on the handle and 60 cm away from the handle (as our navigation targets, particularly for cabinets, do not necessarily allow for optimal viewpoints of the handle). Here as well, we report the best outcome from either navigation targets for RUM.
A11-B Sim2Real Behavior Cloning
Figure A14: Simulated Evaluation of Behavior Cloning. We plot success rate as a function of maximal deviation (meters for translation deviation and radians for rotational deviation). We find that imitation learning particularly struggles to predict accurate robot configurations for cabinet doors.
We compare to a pure imitation learning approach from prior work [21]. More specifically, we train an open-loop policy which, given an RGB-D image of the object after navigation, directly predicts the entire motion plan. As a result, we train on images simulating the robot base around the target navigation locations for cabinets and drawers in the ArtObjSim dataset. The camera pose is fixed to look down at the objects (and slightly to the right for the left-hinged cabinets such that they are in view). As multiple objects may be in view, a red marker is randomly placed somewhere on the object of interest. The target motion plan, consisting of $10$ robot configurations corresponding to the $10$ waypoints in the trajectory, is then mined using SeqIK from the target navigation location.
Once the dataset is generated, we learn a policy to predict the motion plan given an RGB-D image by regressing directly to the joint angles. We experiment with various input modalities (RGB vs RGB-D) and various architectures (fully-connected decoder vs convolutional decoder on top of a pre-trained ResNet encoder) to obtain a policy. We find that passing in depth and using a convolutional decoder helps achieve the lowest validation loss on held-out objects.
We evaluate the performance of our imitation learning model on unseen objects in simulation. An image from the unseen validation set is fed through our policy to predict motion plans. We use forward kinematics to obtain the end-effector pose for each predicted robot configuration, and compute the maximal translational and rotational deviation from the desired end-effector pose. For each object, we record the maximal deviation, which could be translational (in meters) or rotational (in radians), with which we plot success rate as a function of deviation threshold in Figure A14. We compare this to our pipeline in the following manner: we obtain predictions for the articulation parameters using our modified Mask RCNN, and then utilize SeqIK to obtain a motion plan. While our method is only slightly better than imitation learning for drawers, our method significantly outperforms imitation learning on cabinet doors. In particular, the imitation learning model struggles to predict robot configurations with accurate end-effector rotations for the case of cabinet doors.
We evaluate the imitation learning model on a Stretch robot in the real world. One object of each articulation type from the test set is tested. We find that the model struggles to generate robot configurations which are within joint limits when evaluated on real world data. Even after the joint values are clipped to be within joint limits, the robot is unable to even grasp the handle. Our hypothesis is that the sim2real gap is exacerbated by noise in the depth images. Depth images in the real world have random specs and holes close to edges, while the depth images the model is trained on in simulation are a lot cleaner.
A11-C Real World Behavior Cloning
Figure A15: Real World Behavior Cloning. Left. The initial configuration of the robot during data collection and testing. Right. The RGB view from the head camera, with both the end-effector and the drawer visible.
With the sim2real failure of our initial imitation learning baseline, we design an imitation learning baseline in the real world. We adopt a setup similar to Dobb-E [53], except we utilize the head camera on the Stretch as opposed to a wrist camera for policy learning. Here, we collect $32$ demonstrations each for one drawer and one cabinet using tele-operation in the real world. We assume access to our pipeline’s navigation stack, and therefore collect demonstrations by initializing the robot base in a region surrounding the target navigation location. This region is represented by a $15\text{cm}\times 15\text{cm}$ grid with $4\times 4$ base locations, and two demonstrations are collected for each grid point. To ensure consistency between demonstrations, the robot is set to a pre-determined configuration such that the camera can see both the target handle and the end-effector. See Figure A15 for a visualization.
Once the demonstrations have been collected, we learn a closed-loop behavior cloning policy which directly outputs joint angles. We experiment with various input modalities (RGB vs RGB-D), various ways of processing the depth image (average pooling vs ResNet encoder), which time-step of action in the future to predict (e.g. 2nd vs 5th action from the current timestep), and whether initializing the policy with the simulation policy helps or not. The best performing model for the drawer takes in RGB-D, where the depth is processed using average pooling, predicts the 5th action in the future, and is initialized with the simulation policy. As for the cabinet, the best performing model also utilizes RGB-D with average pooling, and is initialized by the simulation policy, but predicts the 10th action in the future. Additionally, we find that filtering out small actions from the dataset helps improve grasping in the real world. This is done by removing samples in each demonstration sequence whose change in joint angles from the previous sample’s do not exceed a specified threshold. Input and label pairings are made using these filtered demonstrations. In summary, both policies leverage ArtObjSim and all real world objects used for development of MOSART (one drawer and two cabinets), making the comparison to MOSART fair.
We deploy our behavior cloning policy on a Stretch robot in the real world. The robot base is initialized to be at an arbitrary location within the $15\text{cm}\times 15\text{cm}$ grid used to collect demonstrations, with the same initial pose and camera orientation. Across $10$ trials, our best model achieves a $70\%$ success rate at opening the seen training drawer used to collect demonstrations. The main failure mode is failure to grasp the handle. We also evaluate our behavior cloning policy on an unseen drawer. Our best performing model is unable to generalize to this novel drawer, achieving a $0\%$ success rate across $10$ trials.
We repeat the above procedure for a left-hinged cabinet. Similar to drawers, our best model is able to open the training cabinet $70\%$ of the time across $10$ trials. Here, the failures were caused from failure to grasp the handle and failure to retract the arm even with an accurate grasp pose. Furthermore, when evaluated on an unseen left-hinged cabinet, the same model achieves a $0\%$ success rate when tested across $10$ trials.
Real world behavior cloning on one object not only fails to generalize to novel objects, but also struggles to consistently open the training object. This result does not come as a complete surprise given that recent works, which do imitation learning in the real world with a similar number of demonstrations [53], do not claim generalization, and need to re-train a separate model for each new drawer or cabinet. It is possible that scaling up imitation learning to a large number of objects may lead to an improvement in performance. However, real world data collection is not easy to scale up. Collecting $32$ tele-operated demonstrations took roughly 2 hours for the drawer, and even longer for the cabinet due to the increased complexity in the motion plan. This makes large-scale imitation learning prohibitively expensive.
A11-D Reinforcement Learning
Prior work has attempted to use end-to-end reinforcement learning to tackle manipulation tasks such as opening cabinets and drawers. ManiSkill2 is a recently proposed benchmark and online challenge for generalizable manipulation skills, and includes the task of interacting with drawers and cabinets. According to the ManiSkill2 Challenge webpage [^1] https://sapien.ucsd.edu/challenges/maniskill/, Gao et al. [17] won the official ManiSkill2 challenge at CVPR 2023’s Embodied AI Workshop, placing first on all tracks. This method uses end-to-end reinforcement learning with the PPO algorithm for rigid-body tasks like interacting with articulated objects. Looking at their results (Table III), we see that on the test set, their method achieves a success rate of 24% on the Open-Cabinet-Door task and 16% and the Open-Cabinet-Drawer task in simulation. A transfer to the real world of this policy has not been demonstrated, and the sim2real gap seems non-trivial – the policy is learned on simulated articulated objects in isolation (ie in the absence of a surrounding environment), unlike articulated objects in the real world, making sim2real transfer very difficult. Furthermore, the success rates of the winning method in simulation are significantly lower than the success rates of our method in the real world.
A11-E Other Prior Systems
Many other past works use privileged information about the object or environment, or design a method for a custom robot. For instance, [56] develops a system for opening various household objects including drawers and cabinet doors, but assumes access to privileged information such as the grasping location, the orientation for the end-effector, and the initial pulling direction. Similarly, [19] presents a framework for opening various kinds of doors, but assumes object detection and utilizes a priori knowledge of the object. [9] develop a graph-search based motion planning algorithm for autonomously opening doors, but also assume access to ground truth articulation parameters. There are many works of this form which make use of privileged information [28, 27, 18]. In summary, all of these works make assumptions which simplifies the problem at hand (e.g. assuming access to the kinematic model of the object), do not provide code, or utilize a custom robot, all of which makes comparison to our zero-shot method on a commodity mobile manipulator difficult.
A12 Comparisons to Prior Work on Articulation Parameter Estimation
Figure A16: Comparison to AO-Grasp. We perform a qualitative comparison to AO-Grasp. AO-Grasp requires a segmented pointcloud – for this, we use APM to segment out objects from our RGB-D images, lift the segmented points to 3D, and then pass this into AO-Grasp to obtain scores for where to grasp. The highest scoring point is treated as the handle location. Our method outperforms AO-Grasp. We believe this is because the handle is not always visible in the noisy point cloud, a drawback of the approach.
Here, we provide more details of our comparison to AO-Grasp [41], as well as a qualitative comparison. AO-Grasp requires a segmented pointcloud; for this, we use APM to segment out objects from our RGB-D images, lift the segmented points to 3D, and then pass this into AO-Grasp to obtain scores for where to grasp. The highest scoring point is treated as the handle location. Figure A16 visualizes a qualitative results. Evidently, the predicted handle is quite off from the ground truth handle in many cases. We believe this is because the handle may not always visible in the noisy point cloud, a drawback of the approach.
A13 Comparison to Direct 3D Prediction
Additionally, we also compare to an approach which directly predicts articulation parameters in 3D. We found that directly predicting the handle in 3D was much worse than our approach (see Table A5, which reports the mean handle error on our real world images). Our hypothesis for the poor performance of direct 3D prediction is that depth in the real world is quite noisy, making the sim2real transfer of a learned neural network model very difficult.
Drawer Left-Hinged Right-Hinged Total
3D Prediction 1.486 m 1.711 m 1.852 m 1.680 m
Ours 0.0317 m 0.0143 m 0.0135 m 0.0211 m
Table A5: Quantitative Comparison to direct 3D prediction. Mean handle error across all instances of our real world testing.